# SMART: Simulated Students Aligned with Item Response Theory for Question Difficulty Prediction

### v1

@@ Abstract
Item (question) difficulties play a crucial role in educational assessments, enabling accurate and efficient assessment of student abilities and personalization to maximize learning outcomes. Traditionally, estimating item difficulties can be costly, requiring real students to respond to items, followed by fitting an item response theory (IRT) model to get item difficulty estimates. This approach cannot be applied to the cold-start setting for previously unseen items either. In this work, we present SMART (Simulated Students Aligned with IRT), a novel method for aligning simulated students with instructed ability, which can then be used in simulations to predict the difficulty of open-ended items. We achieve this alignment using direct preference optimization (DPO), where we form preference pairs based on how likely responses are under a ground-truth IRT model. We perform a simulation by generating thousands of responses, evaluating them with an LLM-based scoring model, and fit the resulting data to an IRT model to obtain item difficulty estimates. Through extensive experiments on a real-world student response dataset, we show that SMART outperforms other item difficulty prediction methods by leveraging its improved ability alignment.
@@ 1 Introduction
Assessing student knowledge and skill levels, an important aspect of education, requires carefully designed assessment items (questions) with high quality and validity. According to the Standards for Educational and Psychological Testing Eignor (2013), various factors play a role in the quality and validity of items, with a crucial factor being their difficulty, i.e., how challenging they are to real students. Item difficulties also play a critical role in learning platforms, enabling personalized learning through student-tailored item recommendation Chen and Duh (2008); Ueno and Miyazawa (2019), item generation Jiao et al. (2023); Kurdi et al. (2020); Shimmei and Matsuda (2022); Ashok Kumar et al. (2023), and curriculum design Mehrens and Phillips (1987); Zhang et al. (2024) to maximize learning outcomes.
@@ 1.1 Contributions
  1. 1.
    First, SMART leverages generative LLMs as the student simulator to synthesize a distribution of student responses to short-answer, open-ended items. These responses provide valuable interpretability in the difficulty estimation process. We use direct preference optimization (DPO) to train our student simulator, using a novel method to form preference pairs. We prefer student responses with higher likelihood under an IRT model, thus aligning the generated responses with both student ability and item difficulty. Second, we use an LLM-based scoring model to score the generated responses using a set of rubrics. Finally, we train an IRT model on the scored student responses to obtain item difficulty estimates.
  2. 2.
    We conduct extensive quantitative experiments on a real-world reading comprehension dataset, with $85$K student responses to $49$ short-answer open-ended items. We find that SMART, by simulating IRT-aligned responses to unseen items from a population of students with different abilities, outperforms state-of-the-art item difficulty prediction methods, even in the low-resource setting with few training items.
  3. 3.
    We analyze the simulated student responses and find that SMART responses correlate strongly with their instructed ability, and that the feature distributions of its responses match closely with the ground-truth. We also qualitatively compare simulated student responses across methods, perform a failure pattern analysis, and suggest directions for future work.
@@ 2 Background: Item Response Theory
$P_{i}(y_{ij}|\theta_{j})=\sigma(\theta_{j}-b_{i}).$ (1)
@@
In our setting, we study short-answer questions where the response is open-ended. We also assume that students can receive partial credit on responses, leading to ordinal scores. A model for ordinal response data is the generalized partial credit model (GPCM) Muraki (1992), which adds a “step” parameter for each of the $C$ total score categories:
$P_{i}(y_{ij}|\theta_{j})=\frac{e^{\sum_{y=0}^{y_{ij}}a_{i}(\theta_{j}-b_{i}+d_{iy})}}{\sum_{c=0}^{C-1}e^{\sum_{y=0}^{c}a_{i}(\theta_{j}-b_{i}+d_{iy})}},$ (2)
where the response $y_{ij}\in\{0,\ldots,C-1\}$, $d_{iy}$ is the step parameter for item $i$ for score $y$, and $a_{i}$ is the discrimination parameter for item $i$. GPCM smoothly adjusts the probability of different ordinal scores according to student ability and item difficulty, with the step parameters determining where the probability mass shifts from one score to the next. Additionally, following prior work Muraki (1993), we set the first step parameter to 00, since only $C-1$ step parameters are required for $C$ score classes. We also impose the location constraint Muraki (1993), where all step parameters for an item sum to 00, i.e., $\sum_{y=0}^{C-1}d_{iy}=0$, ensuring that difficulty parameters across items are on the same scale and are identifiable.
@@ 3 Problem Formulation
We assume that we have access to a real-world dataset of human-scored student responses to open-ended short-answer items. Each item, indexed by $i$, is associated with 1) (optional) metadata $m_{i}$, e.g., a passage text for reading comprehension items, 2) a short question text $q_{i}$, 3) a scoring rubric $w_{i}$, and 4) a training dataset of human-scored student responses denoted by $D_{i}=\{r_{ij},y_{ij}\}$. An IRT model (GPCM) is fit to this data to get ground-truth IRT parameter estimates of the student ability $\theta_{j}$ and item difficulty $b_{i}$. Our goal is to predict the item difficulty parameters of previously unseen, or test items, before students respond to them.
@@ 4 Methodology
Our approach, SMART, illustrated in Figure 1, is a novel, three-stage pipeline, which simulates student responses to short-answer, open-ended items, to help estimate their difficulty. First, SMART leverages human-scored responses by real students to open-ended items to learn a student response simulator, $S(\cdot)$, parameterized by a generative LLM. The simulator $S(\cdot)$ generates a distribution of student responses to unseen open-ended items. Second, an LLM-based scoring model, $R(\cdot)$, scores these generated responses using a set of rubrics. Third, an IRT model, $I(\cdot)$, is trained on the scored student responses to estimate item difficulties.
@@ 4.1 LLM-based Student Response Simulation
We first detail our LLM finetuning method for student simulation, resulting in a model we denote as $S(\cdot)$. We combine the metadata text $m_{i}$, question text $q_{i}$, and the student ability $\theta_{j}$ in our input prompt, and train the LLM using supervised finetuning (SFT) to maximize the likelihood of the ground-truth student response $r_{ij}$ as output. We convert $\theta_{j}$ to a string after rounding to 4 decimal places, which we find works better than binning values into descriptive buckets; we leave an exploration of more advanced methods for numeric inputs to LLMs, such as weighting model parameters Wang et al. (2024), for future work. See the exact prompt format in Appendix E.
@@ 4.2 SMART: Ability Alignment in Response Simulation via Preference Optimization
To address this limitation, we propose a simple idea: to incorporate information on the IRT likelihoods of student responses, in addition to their textual content. The intuition is that generating responses with high likelihoods under the IRT model should lead to better alignment between the IRT parameters and response content, which should benefit the downstream task of item difficulty prediction. To achieve this alignment, we use DPO Rafailov et al. (2023) to prefer responses with scores having a higher IRT likelihood with the input student ability and item difficulty.
@@
Our novel preference pair creation method is illustrated in Figure 2. Given an item’s metadata $m_{i}$, question text $q_{i}$, and a student’s ability $\theta_{j}$, we compare two possible responses: one with a higher IRT likelihood, $r^{w}_{ij}$, which is preferred to another with a lower IRT likelihood $r^{\ell}_{ij}$, thereby forming a preference pair. We find such preference pairs among the human-scored real student responses. For each real student response $r_{ij}$, we construct a set of possible negative candidates, $\mathcal{R}_{ij}$, which are other responses to item $i$ that a student with ability $\theta_{j}$ would be less likely to write, according to the IRT model. Formally:
@@
where $P_{i}$ is the GPCM model from Equation 2, and $\epsilon\in[0,1)$ is a hyperparameter, which sets the threshold for how much more likely a response must be than another to form a preference pair. This threshold is needed to reduce noisy reference pairs. In other words, for some student that responds with $r_{ij}$, for every other possible response we ask the question: how likely is it for the same student to write response $r_{ik}$? We randomly select $m$ negative candidates from $\mathcal{R}_{ij}$ for training, creating $m$ preference pairs where the ground-truth response, $r^{w}_{ij}=r_{ij}$, is preferred over each sampled response, $r^{\ell}_{ij}=r_{ik}$. We note that $\mathcal{R}_{ij}$ could be empty in some cases, for instance, if the ground-truth student response is unlikely for the corresponding $\theta_{j}$.
@@ 4.3 Scoring Model for Open-ended Responses
Given a distribution of varied generated student responses to items, we need to score them before fitting an IRT model (GPCM) to estimate item difficulties. This scoring task is challenging since it requires scoring synthetic responses to items not seen during training. We parameterize our scoring model $R(\cdot)$ by finetuning a generative LLM, given its ability to handle long-context passage inputs, to discern complex (partial) scoring criteria from a rubric and to generalize to scoring responses to new items in the test set. We combine the rubric $w_{i}$, the metadata $m_{i}$, the question text $q_{i}$, and the student response text $r_{ij}$ in a template and instruct the LLM to only output a single integer for the score label $y_{ij}\in\{0,\ldots,C-1\}$. Our scoring model achieves a mean performance of $0.616$ on the quadratic weighted kappa (QWK) metric across $5$ cross-validation folds, which is accurate enough for our purpose; since our goal is not to train the most accurate scoring model, we leave experimenting with other, possibly more advanced scoring models for future work. See the exact prompt we use for our scoring model in Appendix E.
@@ 4.4 Simulation-Based Difficulty Prediction
We now detail how to use our ability-controlled student simulator and automated scoring model to predict item difficulties. First, we create our population of $n$ simulated students such that their abilities match the training distribution. We create a 50-bucket histogram of abilities in the train set, proportionally draw $n$ samples from the resulting distribution, and convert each sample to a random uniform value in the corresponding bin. Second, we perform the student simulation. For each drawn ability and item in the test set, we prompt the simulated student model $S(\cdot)$ to generate a corresponding response using nucleus sampling. Third, we assign the most likely score to each generated response using the automated scoring model $R(\cdot)$. Finally, we re-fit the IRT model $I(\cdot)$ using the new set of scored responses. For each generated response, we assign a student ID rounding the ability to the nearest decimal place, which we find to be crucial in preventing overfitting by reducing the number of ability parameters significantly. In this step, we also include the ground truth student responses from the training set and fit the IRT model on both real and generated responses simultaneously, to keep the IRT model close to its calibrated version fitted on the training set. We take the difficulty estimates under this IRT model for the new, unseen items as our final predictions.
@@ 5 Experimental Evaluation
We now detail experiments that we conduct to validate our approach on real student response data.
5.1 Dataset
We conduct our experiments on a large real-world dataset of student responses to open-ended short answer items, administered to tens of thousands of students in the United States, containing around $85$K responses to $49$ items from around $63$K students. These open-ended items assess reading comprehension skills including argumentation, explanation, and narration, at the grade $6$ level with the average length of student responses being $92$ tokens. Each response is manually scored by a trained human scorer, using a rubric with detailed scoring criteria, with scores in the set $\{0,1,2\}$. We show an example item in Table 2. We work in a low resource setting with few training items, which is particularly challenging for existing state-of-the-art difficulty prediction methods AlKhuzaey et al. (2024); Benedetto (2023); Benedetto et al. (2023), which finetune encoder-based LLMs that rely on a large number of training items.
We split the dataset into $5$ cross-validation folds, each with $29$/$10$/$10$ items in train/validation/test, respectively. We fit IRT models (GPCM) on each fold to estimate ground truth student ability and item parameters. For each fold, to ensure no information from unseen items is leaked at train time, we fit three separate IRT models, one on the train data, one on train plus validation data, and one on all the data. We use the estimated parameters from the three models to train our student simulator, to perform validation, and to perform test-time evaluation, respectively. We ensure that the distribution of difficulties of items across all splits and folds is roughly equal. We provide further details on dataset processing in Appendix B.
@@ 5.2 Metrics
Item Difficulty Prediction
Following prior work in item difficulty prediction Huang et al. (2017a); He-Yueya et al. (2024), we use the widely adopted Pearson Correlation Coefficient (PCC) to measure the linear correlation between predicted and ground-truth item difficulties, Spearman Rank Correlation Coefficient (SCC) to measure the correlation between ranks of predicted and ground-truth item difficulties, and the conventional Root Mean Squared Error (RMSE) for a comparison between predicted and ground-truth item difficulties.
Student Simulation
We additionally propose metrics to investigate the properties of the generated responses. Since we are modeling students at a population level, reference-based metrics such as ROUGE-L Lin (2004) and BERTScore Zhang et al. (2020) are unsuitable to evaluate response quality. Therefore, following prior work in population simulation Yu et al. (2023); Bui et al. (2025), we report distribution-level metrics comparing the generated and ground-truth response distributions in a latent embedding space. We first extract embeddings for all responses using a sentence encoder Reimers and Gurevych (2019), specifically multi-qa-mpnet-base-dot-v1. We then compute the Fréchet Inception Distance (FIDHeusel et al. (2017), which is common for evaluating generated images but has recently been adapted to text Bui et al. (2025); Yue et al. (2023), and MAUVE Pillutla et al. (2021), which measures the similarity of cluster distributions in the embedding space, to evaluate the alignment between the predicted and ground-truth student response distributions. Following Bui et al. (2025), we also measure the similarity in diversity distribution by computing Diversity KL, the KL divergence between the histograms of pairwise response cosine similarity across all ground-truth responses and across all simulated responses. Finally, we compute $\theta$-Align, the Spearman correlation between input student ability and the predicted score of the corresponding generated response, measuring how well simulated students align with their prompted abilities.

### v2

@@ Abstract
Item (question) difficulties play a crucial role in educational assessments, enabling accurate and efficient assessment of student abilities and personalization to maximize learning outcomes. Traditionally, estimating item difficulties can be costly, requiring real students to respond to items, followed by fitting an item response theory (IRT) model to get difficulty estimates. This approach cannot be applied to the cold-start setting for previously unseen items either. In this work, we present SMART (Simulated Students Aligned with IRT), a novel method for aligning simulated students with instructed ability, which can then be used in simulations to predict the difficulty of open-ended items. We achieve this alignment using direct preference optimization (DPO), where we form preference pairs based on how likely responses are under a ground-truth IRT model. We perform a simulation by generating thousands of responses, evaluating them with a large language model (LLM)-based scoring model, and fit the resulting data to an IRT model to obtain item difficulty estimates. Through extensive experiments on two real-world student response datasets, we show that SMART outperforms other item difficulty prediction methods by leveraging its improved ability alignment.
@@ 1 Introduction
Assessing student knowledge and skill levels, an important aspect of education, requires carefully designed assessment items (i.e., questions) with high quality and validity. According to the Standards for Educational and Psychological Testing Eignor (2013), various factors play a role in the quality and validity of items, with a crucial factor being their difficulty, i.e., how challenging they are to real students. Item difficulties also play a critical role in learning platforms, enabling personalized learning through student-tailored item recommendation Chen and Duh (2008); Ueno and Miyazawa (2019), item generation Jiao et al. (2023); Kurdi et al. (2020); Shimmei and Matsuda (2022); Ashok Kumar et al. (2023), and curriculum design Mehrens and Phillips (1987); Zhang et al. (2024) to maximize learning outcomes.
@@ 1.1 Contributions
First, SMART leverages generative LLMs as the student simulator to synthesize a distribution of student responses to open-ended items. These responses provide valuable interpretability in the difficulty estimation process. We use direct preference optimization (DPO) to train our student simulator, using a novel method to form preference pairs. We prefer student responses with higher likelihood under an IRT model, thus aligning the generated responses with both student ability and item difficulty. Second, we use an LLM-based scoring model to score the generated responses. Finally, we train an IRT model on the scored student responses to obtain item difficulty estimates.
We conduct extensive quantitative experiments on two real-world datasets, one with 49 short-answer reading comprehension items and another with 50 Java coding items. We find that SMART, by simulating IRT-aligned responses to unseen items from a population of students with different abilities, outperforms state-of-the-art item difficulty prediction methods, even in the low-resource setting with few training items. We analyze the simulated student responses and find that SMART responses correlate strongly with their instructed ability, and that the feature distributions of its responses match closely with the ground-truth. We also qualitatively compare simulated student responses across methods, perform a failure pattern analysis, and suggest directions for future work.
@@ 2 Background: Item Response Theory
$P^{\text{1PL}}_{i}(y_{ij}|\theta_{j})=\sigma(\theta_{j}-b_{i}).$ (1)
@@
In our setting, we study short-answer questions where the response is open-ended. We assume that students can receive partial credit on responses, leading to ordinal or continuous scores. For ordinal response data, we use the generalized partial credit model (GPCM) Muraki (1992), which adds a “step” parameter for each of the $C$ total score categories:
$P^{\text{GPCM}}_{i}(y_{ij}|\theta_{j})=\frac{e^{\sum_{y=0}^{y_{ij}}a_{i}(\theta_{j}-b_{i}+d_{iy})}}{\sum_{c=0}^{C-1}e^{\sum_{y=0}^{c}a_{i}(\theta_{j}-b_{i}+d_{iy})}},$ (2)
where the response $y_{ij}\in\{0,\ldots,C-1\}$, $d_{iy}$ is the step parameter for item $i$ for score $y$, and $a_{i}$ is the discrimination parameter for item $i$. GPCM smoothly adjusts the probability of different ordinal scores according to student ability and item difficulty, with step parameters determining where the probability mass shifts from one score to the next.
We also handle continuous scores, i.e., $y_{ij}\in[0,1]$, where the student’s response lies on a spectrum between incorrect and correct. In this setting, we model correctness probability using a continuous Bernoulli distribution Loaiza-Ganem and Cunningham (2019):
$P^{\text{Bern.}}_{i}(y_{ij}|\theta_{j})=$ $P^{\text{1PL}}_{i}(Y=1|\theta_{j})^{y_{ij}}\cdot$ (3)
$P^{\text{1PL}}_{i}(Y=0|\theta_{j})^{(1-y_{ij})}.$
@@ 3 Problem Formulation
We assume that we have access to a real-world dataset of human-scored student responses to open-ended short-answer items. Each item, indexed by $i$, is associated with 1) a short question text $q_{i}$, 2) (optional) meta data $m_{i}$, e.g., a passage text for reading comprehension items, 3) (optional) a scoring rubric $w_{i}$, and 4) a set of human-scored student responses denoted by $D_{i}=\{r_{ij},y_{ij}\}$. An IRT model is fit to this data to get ground-truth estimates of each student ability $\theta_{j}$ and item difficulty $b_{i}$. Our goal is to predict the item difficulty parameters of previously unseen test items before students respond to them.
@@ 4 Methodology
Our approach, SMART, illustrated in Figure 1, uses a novel, three-stage pipeline, which simulates student responses to short-answer, open-ended items, to help estimate their difficulty. First, SMART leverages human-scored responses by real students to open-ended items to learn a student response simulator, $S(\cdot)$, parameterized by a generative LLM. The simulator $S(\cdot)$ generates a distribution of student responses to unseen open-ended items. Second, an LLM-based scoring model, $R(\cdot)$, scores these generated responses. Third, an IRT model, $I(\cdot)$, is trained on the scored student responses to estimate item difficulties.
@@ 4.1 LLM-based Student Response Simulation
We first detail our LLM finetuning method for student simulation, resulting in a model we denote as $S(\cdot)$. We combine the question text $q_{i}$, the meta data text $m_{i}$, and the student ability $\theta_{j}$ in our input prompt, and train the LLM using supervised finetuning (SFT) to maximize the likelihood of the ground-truth student response $r_{ij}$ as output. We convert $\theta_{j}$ to a string after rounding to 4 decimal places, which we find works better than binning values into descriptive buckets; we leave an exploration of more advanced methods for numeric inputs to LLMs, such as weighting model parameters Wang et al. (2024), for future work. See the exact prompt format in Appendix F.
@@ 4.2 SMART: Ability Alignment in Response Simulation via Preference Optimization
To address this limitation, we propose the following idea: to incorporate information on the IRT likelihoods of student responses, in addition to their textual content. The intuition is that generating responses with high likelihoods under the IRT model should lead to better alignment between the IRT parameters and response content, which should benefit the downstream task of item difficulty prediction. To achieve this alignment, we use DPO Rafailov et al. (2023) to prefer responses with scores having a higher IRT likelihood with the input student ability and item difficulty.
@@
Our novel preference pair creation method is illustrated in Figure 2. Given an item’s question text $q_{i}$, meta data $m_{i}$, and a student’s ability $\theta_{j}$, we compare two possible responses: one with a higher IRT likelihood, $r^{w}_{ij}$, which is preferred to another with a lower IRT likelihood $r^{\ell}_{ij}$, thereby forming a preference pair. We find such preference pairs among the human-scored real student responses. For each real student response $r_{ij}$, we construct a set of possible negative candidates, $\mathcal{R}_{ij}$, which are other responses to item $i$ that a student with ability $\theta_{j}$ would be less likely to write, according to the IRT model. Formally:
@@
where $P_{i}$ is either $P^{\text{GPCM}}_{i}$ (Eq. 2) if the scores are ordinal or $P^{\text{Bern.}}_{i}$ (Eq. 3) if the scores are continuous. The hyperparameter $\epsilon\in[0,1)$ sets the threshold for how much more likely a response must be than another to form a preference pair, which is needed to reduce noisy preference pairs. In summary, for some student that responds with $r_{ij}$, for every other possible response, we ask the question: how likely is it for the same student to write response $r_{ik}$? We randomly select $m$ negative candidates from $\mathcal{R}_{ij}$ for training, creating $m$ preference pairs where the ground-truth response, $r^{w}_{ij}=r_{ij}$, is preferred over each sampled response, $r^{\ell}_{ij}=r_{ik}$. We note that $\mathcal{R}_{ij}$ could be empty in some cases, for instance, if the ground-truth student response is unlikely for the corresponding $\theta_{j}$.
@@ 4.3 Scoring Model for Open-ended Responses
Given a distribution of varied generated student responses to items, we need to score them before fitting an IRT model to estimate item difficulties. This scoring task is challenging since it requires scoring synthetic responses to items not seen during training. We parameterize our scoring model $R(\cdot)$ by finetuning a generative LLM, given its ability to handle long-context inputs, to discern complex (partial) scoring criteria, optionally from a rubric, and to generalize to scoring responses to new items in the test set. We combine the question text $q_{i}$, the meta data $m_{i}$, the scoring rubric $w_{i}$, and the student response text $r_{ij}$ into a prompt $x_{ij}$, and instruct the LLM to only output a single integer for the score label $y_{ij}\in\{0,\ldots,C-1\}$. For ordinal scores, we produce an estimate score $\hat{y}_{ij}$ via greedy decoding from $R(\cdot|x_{ij})$. For continuous scores, following Scarlatos et al. (2025a), we smoothly interpolate between the token logits corresponding to fully correct and incorrect scores:
$\hat{y}_{ij}=\frac{e^{R(1|x_{ij})}}{e^{R(0|x_{ij})}+e^{R(1|x_{ij})}},$ (5)
where $R(y|x_{ij})$ is the output logit value for token $y$ returned by $R(\cdot|x_{ij})$, and we train using binary cross-entropy loss. See the exact prompt we use for our scoring model in Appendix F.
@@ 4.4 Simulation-Based Difficulty Prediction
We now detail how to use our ability-controlled student simulator and automated scoring model to predict item difficulties. First, we create a population of $n$ simulated students such that their abilities match the training distribution: we create a 50-bucket histogram of abilities in the train set, proportionally draw $n$ samples from the resulting distribution, and convert each sample to a random uniform value in the corresponding bin. Second, we perform the student simulation: for each drawn ability and item in the test set, we prompt the simulated student model $S(\cdot)$ to generate a corresponding response using nucleus sampling. Third, we score each generated response using the automated scoring model $R(\cdot)$. Finally, we re-fit the IRT model $I(\cdot)$ using the new set of scored responses. For each generated response, we assign a student ID by rounding the ability to the nearest decimal place, which we find to be crucial in preventing overfitting by reducing the number of ability parameters significantly. We provide additional details on IRT fitting in Appendix A.4. We take the difficulty estimates under this IRT model for the new, unseen items as our final predictions.
@@ 5 Experimental Evaluation
We now detail experiments that we conduct to validate our approach on real student response datasets.
5.1 Datasets
We conduct our experiments on two real-world open-ended student response datasets, each containing around 50 items and thousands of student responses.
Smarter Balanced.
We first use a private real-world dataset, provided to us by Smarter Balanced [^1] https://smarterbalanced.org/. The dataset contains student responses to open-ended short answer items administered to students in the United States, containing around 8585K responses to $49$ items from around 6363K students. These open-ended items assess reading comprehension skills including argumentation, explanation, and narration, at the grade $6$ level with the average length of student responses being $92$ tokens. Each response is manually scored by a trained human scorer, using a rubric with detailed scoring criteria, with scores in the set $\{0,1,2\}$. The meta data for each item includes a passage that the student must read before answering the question. Our trained 3B parameter scoring model achieves a quadratic weighted kappa (QWK) of $0.616$ on this dataset, indicating moderate-to-high agreement with ground-truth scores, which is competitive with unseen item open-ended response scoring on other datasets [^2] Best QWK on unseen items in NAEP Reading Automated Scoring Challenge was $0.528$: https://github.com/NAEP-AS-Challenge/reading-prediction/blob/main/results.md. We show example items in Tables 3 and 4.
CodeWorkout.
We additionally use the publicly available CodeWorkout dataset DataShop (2021), originally used for the 2nd CSEDM Data Challenge Organizers (2021). The dataset, commonly used for student modeling in open-ended programming tasks Liu et al. (2022); Duan et al. (2025), contains 50 Java coding problems in a first year college programming course, with student solutions across 2 semesters. Students can attempt a problem multiple times after receiving compiler feedback, so we only consider the first attempt for a problem for each student, resulting in 10,834 responses from 246 students. Each response is scored with a continuous value between 0 and 1, representing the portion of test cases that passed; the test cases themselves are not made available. To ensure that models understand how to correctly solve the problems, we include 2 LLM-generated solutions in each item’s meta data, using the method described in Duan et al. (2025). Our trained 3B parameter scoring model achieves a Pearson’s correlation coefficient (PCC) of $0.849$ on this dataset, indicating high agreement with ground-truth scores. We show an example item in Table LABEL:tab:qual_ex_codeworkout.
We split each dataset into $5$ cross-validation folds, with each fold split 60/20/20 item-wise for train/validation/test, respectively. We ensure that difficulties of items across all splits and folds are roughly even. We fit IRT models ($P^{\text{GPCM}}$ for Smarter Balanced and $P^{\text{Bern.}}$ for CodeWorkout) on each fold to estimate ground-truth student ability and item parameters. For each fold, to ensure no information from unseen items is leaked at train time, we fit three separate IRT models, one on train data, one on train plus validation data, and one on all data. We use the estimated parameters from the three models to train our student simulator, and to perform validation and test-time evaluation, respectively. We provide further details on IRT fitting in Appendix A.4 and dataset processing in Appendix B.
@@ 5.2 Metrics
Item Difficulty Prediction.
Following prior work in item difficulty prediction Huang et al. (2017a); He-Yueya et al. (2024), we use the widely adopted Pearson Correlation Coefficient (PCC) to measure the linear correlation between predicted and ground-truth item difficulties, Spearman Rank Correlation Coefficient (SCC) to measure the correlation between ranks of predicted and ground-truth item difficulties, and Root Mean Squared Error (RMSE) for a precision comparison between predicted and ground-truth item difficulties.
Student Simulation.
We additionally propose metrics to investigate how faithful the generated responses are to real student ones. Since we model students at a population level, reference-based metrics such as ROUGE-L Lin (2004) and BERTScore Zhang et al. (2020) are not suitable. Therefore, following prior work in population simulation Bui et al. (2025); Yu et al. (2023), we report distribution-level metrics comparing the generated and ground-truth response distributions in a latent embedding space. We first extract embeddings for all responses using a sentence encoder Reimers and Gurevych (2019), specifically multi-qa-mpnet-base-dot-v1. We then compute the Fréchet Inception Distance (FIDHeusel et al. (2017), which is common for evaluating generated images but has recently been adapted to text Yue et al. (2023); Bui et al. (2025), and MAUVE Pillutla et al. (2021), which measures the similarity of cluster distributions in the embedding space, to evaluate the alignment between the predicted and ground-truth student response distributions. Following Bui et al. (2025), we also measure the similarity in diversity distribution by computing Div. KL, the KL divergence between the histograms of pairwise response cosine similarity across all ground-truth responses and across all simulated responses. Finally, we compute $\theta$-Align, the Spearman correlation between input student ability and the predicted score of the corresponding generated response, measuring how well simulated students align with their prompted abilities.